d²z/dx²=sin y, solve it by 2 order integration
1
Expert's answer
2020-11-18T19:28:51-0500

Integrate with respect x


"\frac{dz}{dx}=siny \times x + c_1"


"dz = siny \times xdx +c_1dx"


Again integrate


"z = siny\frac{x^2}{2} +c_1x+c"